# Can a Multichoice Dataset be Repurposed for Extractive Question Answering?

### v1

@@ 4.6 Inter-Annotator Agreement
We carried out an inter-annotation (IAA) study in order to assess the level of agreement across annotators, and get an indication of how reliable the guidelines were. [^3] As the dataset was small and fully manually reviewed, the IAA study did not serve to evaluate the data quality. For the IAA study, we selected 50 English QA pairs for annotation by three separate annotators. The three annotations for each QA pair were then compared.
Instead of calculating an F-score measure, which would give indication of quality of annotations, we utilize the $\gamma$ measure for Inter-Annotator Agreement and Alignment Mathet et al. (2015). [^4] Gamma ranges between 0 and 1 where a higher value indicates higher agreement between annotators. This measure involves identifying the optimal alignment among annotations provided by multiple annotators, aiming to align annotations with high similarity and compute the average dissimilarity between them. Specifically, the average dissimilarity is calculated by averaging dissimilarities among aligned annotations, considering both their categories and positions. The alignment with the minimum mean dissimilarity, selected from all potential alignments initially evaluated during computation, is employed. The final inter-annotator agreement ($\gamma$ measure) was computed on the three sets of 50 QA pairs yielding a high $\gamma$ value of $0.81$, indicating that the designed annotation guidelines are reliable. Given that the MSA annotation guidelines were based on the EN guidelines, we posit that their quality was just as high. The three sets of 50 EN IAA QAs were merged and reviewed to produce a final version for inclusion in our parallel dataset. See Appendix A for more details and visualization of the the process of combining the spans.
@@ 5.1 Experimental Setup
5.1.1 Datasets
We use two datasets in our experiments: the commonly used state-of-the art QA dataset, SQuAD Rajpurkar et al. (2016) and our Belebele-EQA, both of which are Wikipedia-based. For SQuAD, we only used the validation portion of the dataset which contains 10,657 English EQA pairs. For Belebele-EQA, we have two versions: (i) Belebele-EQA-All contains 329 QA pairs for EN, MSA and five dialects, and (ii) Belebele-EQA-Sub excludes all BB-annotated questions, leaving 284 QA pairs for English and 297 for MSA. The difference in Sub numbers reflects the difficulty of the task of labelling QAs for their level of complexity due to linguistic differences and annotator interpretation.
5.1.2 QA Evaluation Tool
To run the QA task on the above datasets, we used PrimeQA Sil et al. (2023), an open-source [^5] https://github.com/primeqa/primeqa tool that is built on top of the HuggingFace Wolf et al. (2020) library to support research on running different QA tasks. We used two existing models, namely PrimeQA/NQ [^6] PrimeQA/nq_tydi-reader-xlmr_large-20221210 and PrimeQA/NQ+SQuAD, [^7] PrimeQA/nq_tydi_sq1-reader-xlmr_large-20221110 which are 550m parameters XLM-RLarge Conneau et al. (2020) based models that were finetuned for the multilingual TyDi QA task Clark et al. (2020). Both models were finetuned on the Natural Questions dataset Kwiatkowski et al. (2019). The latter, however, was further fine-tuned on the training portion of the SQuAD dataset.
5.1.3 Evaluation Metrics
@@ 5.2 Results and Analysis
5.2.1 SQuAD vs. BELEBELE-EQA
We evaluate the SQuAD dataset as well as the EN-EN QA pairs from the Belebele-EQA dataset using the PrimeQA toolkit. This experiment allows us to (a) verify that the chosen pretrained models are capable of solving the QA task at-hand without further finetuning, and (b) establish a sense of difficulty for the Belebele-EQA datasets compared to the commonly used SQuAD. We observe the following based on the results in Table 2: First, on SQuAD, the F1 shows a negligible difference between the two pretrained models, while EM is surprisingly lower by 3.4% for the model that was finetuned on the SQuAD dataset. [^8] The state-of-the-art F1 and EM scores on SQuAD are 93.2 and 90.9, respectively, as per the public PapersWithCode leaderboard: https://paperswithcode.com/sota/question-answering-on-squad20
In contrast, the results on Belebele-EQA-All show that additional finetuning on SQuAD yielded better performance both in F1 and EM.
Second, the best Belebele-EQA (All and Sub) F1 scores (71.0 and 76.4) are 18.6 and 14.2 points less than SQuAD’s F1 score, respectively. Similarly, the best respective EM scores are 50.5 and 56, or 28.6 and 23.1 points less than that of the SQuAD dataset. Finally, and as expected, removing the Belebele-EQA hard questions (Belebele-EQA-Sub) resulted in better performance on both F1 and EM measures.
5.2.2 BELEBELE-EQA Results
Based on the baseline results presented above, we report in Table 3 on the different question language setups using the best identified model: PrimeQA/NQ+SQuAD. For specific experiments, we refer to the pair of languages in passage-question, e.g., EN-EN, MSA-Iraqi, or MSA-DA.
@@ 5.2.2 BELEBELE-EQA Results
Belebele-EQA-Sub 284 71.6 51.0
PrimeQA/NQ+SQuAD
@@
DA–Levantine 49.5 31.9 51.2 36.5 52.3 35.2 54.0 39.8
DA–Gulf 50.4 31.3 52.6 37.4 52.4 33.1 54.5 39.1
DA–Moroccan 36.2 18.5 38.3 21.0 37.0 20.1 39.2 22.2
DA–Egyptian 48.7 30.1 51.2 36.2 51.4 33.5 54.0 39.4
@@
DA–Levantine 54.5 35.6 56.0 38.9 56.5 37.0 58.2 40.7
DA–Gulf 56.6 36.2 58.4 40.7 57.2 36.7 59.2 41.8
DA–Moroccan 44.5 26.1 46.0 29.5 46.4 27.6 48.1 31.3
DA–Egyptian 54.8 34.3 56.9 39.5 55.9 35.4 58.2 40.7
@@
Table 3: Monolingual and Cross-lingual evaluation of the Belebele-EQA dataset using the PrimeQA toolkit with PrimeQA/NQ+SQuAD. (EM: Exact Match; F1n and EMn are the text-normalized versions of F1 and EM, respectively; DA: Dialectal Arabic; Scores are percentages; Higher is better.)
Cross-Lingual Experiments (MSA-EN, EN-MSA
@@
The DA questions consistently lead to lower performance compare to the above-mentioned settings, and with performance with MSA passages being better on average than on EN passages (by 6.3% absolute). The specific dialects seem to follow a common order of Gulf$>$(Egyptian|Levantine)$>$Iraqi$>$Moroccan, which we speculate may reflect the order of their question sets’ similarity to MSA. An investigation in the differences in lengths between references and predictions shows that the DA question systems had an average of 2.1 extra words per prediction in EN (over 3.5 words in reference), and 0.5 extra words in MSA (over 3.4 words in reference). This reference-prediction length difference had a correlation of 79.1% with the systems F1-score.
@@ 6 Discussion
Answering the research question “Can a Multichoice Dataset be Repurposed for Extractive Question Answering?” involves considering the various insights and findings presented in this paper. Regarding the use of Belebele, we found that repurposing is possible, resulting in the creation of new datasets and benchmarks. However, our exploratory study required significant manual effort. Contrary to initial assumptions, we discovered that the original answers from the multiple-choice question pairs were not suitable for automated or manual approaches.
We can see that the repurposing process is much more feasible when the differences between the QA styles are fully understood. We found that, through an effective filtering process, we could easily eliminate many QA pairs that did not lend themselves easily to EQA. While this resulted in reducing the dataset size by more than half, this leaves open new avenues for examining how those ‘unsuitable’ QA pairs could be reformulated to be useful in EQA. In addition, the dataset style differences explained in Table 1 create challenges for direct mapping from question to answer span. Our manual efforts and the subsequent insights can now serve to make this repurposing task much more feasible and potentially automated by other dataset developers.
Repurposing a multilingual dataset like Belebele means that its parallelism could be exploited through a bootstrapping approach. We have learned that once one language dataset is manually annotated and verified, the spans could be relatively easily projected to a new language from the wider dataset provided a translation system is available. Likewise, few changes were required when adapting our annotation guidelines from EN to MSA, suggesting that this adaptability may hold for other languages.
Finally, our experimental results confirm that our approach produced a useful data set that can be used for benchmarking cross-lingual EQA for for all Belebele languages.
@@ 7 Conclusion and Future Work
We repurposed the MCQA Belebele dataset to create a new EQA dataset. Our efforts yielded a new extractive QA dataset for English and MSA, accompanied by cross-lingual benchmark results for dialectal Arabic QA.999 The dataset will be made publicly available. We also provided detailed guidelines beneficial for practitioners exploring question-answering in other languages within the Belebele dataset. Our work aims to enhance comprehension of the challenges and opportunities in reformulating tasks within NLP research.
For future work, we would like to support the repurposing of Belebele-EQA for other languages based on our experiences and insights, with the potential of automating the process. We are also interested in using the dataset as a basis for fine-tuning models for general purpose Arabic QA. Finally, based on the insights we gained from our error analysis, we aim to examine how to improve EQA evaluation metrics, with specific consideration of the role of span length.
@@
We would like to point to some limitations of our study. First, while we are addressing a general problem, in our study, out of all the other Belebele under-resourced languages, we only focused on cross-lingual evaluation of five Arabic dialects. While we believe that our findings are more generally applicable, we would like to note that each language and dataset may present unique challenges and opportunities. Second, our repurposing was based on a single dataset, Belebele, and for a single task (extractive question answering in the style of machine reading comprehension), and we note that other tasks may require some specific considerations. Finally, our repurposing is limited to languages for which there exist datasets for the original task.
@@
It is important to acknowledge the possibility of biases within our dataset due to subjective human judgments, or due to the influence of the original task’s data.
@@ Appendix A Processing IAA Spans for Visualization
Appendix B Arabic Examples of MCQ-Span Types
@@
Appendix C Guidelines for English Answer Span Annotation
Appendix D Guidelines for Modern Standard Arabic Answer Span Annotation

### v2

@@ 4.6 Inter-Annotator Agreement
We carried out an inter-annotation (IAA) study in order to assess the level of agreement across annotators, and get an indication of how reliable the guidelines were. [^3] As the dataset was small and fully manually reviewed, the IAA study did not serve to evaluate the data quality. For the IAA study, we selected 50 English QA pairs for annotation by three separate annotators. The three annotations for each QA pair were then compared. Instead of calculating an F-score measure, which would give indication of quality of annotations, we utilize the $\gamma$ measure for Inter-Annotator Agreement and Alignment Mathet et al. (2015). [^4] Gamma ranges between 0 and 1 where a higher value indicates higher agreement between annotators. This measure involves identifying the optimal alignment among annotations provided by multiple annotators, aiming to align annotations with high similarity and compute the average dissimilarity between them. Specifically, the average dissimilarity is calculated by averaging dissimilarities among aligned annotations, considering both their categories and positions. The alignment with the minimum mean dissimilarity, selected from all potential alignments initially evaluated during computation, is employed. The final inter-annotator agreement ($\gamma$ measure) was computed on the three sets of 50 QA pairs yielding a high $\gamma$ value of $0.81$, indicating that the designed annotation guidelines are reliable. Given that the MSA annotation guidelines were based on the EN guidelines, we posit that their quality was just as high. The three sets of 50 EN IAA QAs were merged and reviewed to produce a final version for inclusion in our parallel dataset. See Appendix A for more details and visualization of the span combination process.
@@ 5.1 Experimental Setup
Datasets
We use two datasets in our experiments: the commonly used state-of-the art QA dataset, SQuAD Rajpurkar et al. (2016) and our Belebele-EQA, both of which are Wikipedia-based. For SQuAD, we only used the validation portion of the dataset which contains 10,657 English EQA pairs. For Belebele-EQA, we have two versions: (i) Belebele-EQA-All contains 329 QA pairs for EN, MSA and five dialects, and (ii) Belebele-EQA-Sub excludes 44 BB-annotated questions, leaving 285 QA pairs.
QA Evaluation Tool
To run the QA task on the above datasets, we used PrimeQA Sil et al. (2023), an open-source [^5] https://github.com/primeqa/primeqa tool that is built on top of the HuggingFace Wolf et al. (2020) library to support research on running different QA tasks. We used two existing models, namely PrimeQA/NQ+TyDi [^6] PrimeQA/nq_tydi-reader-xlmr_large-20221210 and PrimeQA/NQ+TyDi+SQuAD, [^7] PrimeQA/nq_tydi_sq1-reader-xlmr_large-20221110 which are 550m parameters XLM-RLarge Conneau et al. (2020) based models that were finetuned for the multilingual TyDi QA task Clark et al. (2020). Both models were finetuned on the Natural Questions dataset Kwiatkowski et al. (2019) and TyDi Clark et al. (2020). The latter, however, was further fine-tuned on the training portion of the SQuAD dataset.
Evaluation Metrics
@@ 5.2 Results and Analysis
SQuAD vs. BELEBELE-EQA
We evaluate the SQuAD dataset as well as the EN-EN QA pairs from the Belebele-EQA dataset using the PrimeQA toolkit (See Appendix B for details) This experiment allows us to (a) verify that the chosen pretrained models are capable of solving the QA task at-hand without further finetuning, and (b) establish a sense of difficulty for the Belebele-EQA datasets compared to the commonly used SQuAD. We observe the following based on the results in Table 2: First, on SQuAD, the F1 shows a negligible difference between the two pretrained models, while EM is surprisingly lower by 3.4% for the model that was finetuned on the SQuAD dataset. [^8] The state-of-the-art F1 and EM scores on SQuAD are 93.2 and 90.9, respectively, as per the public PapersWithCode leaderboard: https://paperswithcode.com/sota/question-answering-on-squad20 In contrast, the results on Belebele-EQA-All show that additional finetuning on SQuAD yielded better performance both in F1 and EM.
Second, the best Belebele-EQA (All and Sub) F1 scores (71.0 and 76.6) are 18.6 and 14 points less than SQuAD’s F1 score, respectively. Similarly, the best respective EM scores are 50.5 and 56.1, or 28.6 and 23 points less than that of the SQuAD dataset. Finally, and as expected, removing the Belebele-EQA hard questions (Belebele-EQA-Sub) resulted in better performance on both F1 and EM measures.
BELEBELE-EQA Results
Based on the baseline results presented above, we report in Table 3 on the different question language setups using the best identified model: PrimeQA/NQ+TyDi+SQuAD. For specific experiments, we refer to the pair of languages in passage-question, e.g., EN-EN, or MSA-Iraqi.
@@
Belebele-EQA-Sub 285 71.3 50.5
PrimeQA/NQ+TyDi+SQuAD
@@
DA–Levantine 49.5 31.9 51.2 36.5 51.9 35.1 53.7 39.6
DA–Gulf 50.4 31.3 52.6 37.4 52.1 33.0 54.2 38.9
DA–Morrocan 36.3 18.5 38.5 21.0 37.0 20.0 39.3 22.1
DA–Egyptian 48.8 30.1 51.5 36.2 51.3 33.3 54.0 39.3
DA–Average 45.9 27.9 48.0 32.5 47.7 30.2 49.8 34.6
@@
DA–Levantine 54.2 35.3 55.7 38.6 58.5 39.3 60.4 43.2
DA–Gulf 56.6 36.2 58.4 40.7 59.3 39.6 61.3 44.2
DA–Morrocan 45.1 26.7 46.6 30.1 48.1 29.8 49.6 33.3
DA–Egyptian 54.5 34.0 56.6 39.2 57.3 37.5 59.5 42.5
DA–Average 52.1 32.9 53.7 36.7 55.5 36.5 57.3 40.5
@@
DA–Levantine (TMSA) 58.5 38.3 60.3 43.2 62.1 42.1 64.1 46.7
DA–Gulf (TMSA) 57.9 38.0 59.9 42.9 61.0 40.7 63.3 46.0
DA–Morrocan (TMSA) 56.5 36.2 58.6 42.6 59.4 39.3 61.5 45.3
DA–Egyptian (TMSA) 58.5 37.7 60.4 42.2 62.1 41.1 64.2 46.0
DA–Average (TMSA) 56.6 36.3 58.5 41.2 59.7 39.4 61.8 44.4
Table 3: Monolingual and Cross-lingual evaluation of the Belebele-EQA dataset using the PrimeQA toolkit with PrimeQA/NQ+TyDi+SQuAD. (EM: Exact Match; F1n and EMn are the text-normalized versions of F1 and EM, respectively; DA: Dialectal Arabic; TEN/MSA: Question translated to EN/MSA; Scores are %s; Higher is better.)
Cross-Lingual Experiments (MSA-EN, EN-MSA)
@@
The DA questions consistently lead to lower performance compare to the above-mentioned settings, and with performance with MSA passages being better on average than on EN passages (by 6.3% absolute). The specific dialects seem to follow a common order of Gulf$>$(Egyptian|Levantine)$>$Iraqi$>$Moroccan, which we speculate may reflect the order of their question sets’ similarity to MSA. Automatically translating dialectal questions into the target passage language using Google Translate [^9] For DA-to-English, we translated using Google Translate (AR-to-EN), while for DA-to-MSA, we used English as an intermediate language (DA $\Rightarrow$ EN $\Rightarrow$ MSA): Google Translate (AR-to-EN $\Rightarrow$ EN-to-AR). results in notable improvements: an average increase of 18% for English passages and 4.4% for MSA passages (All setting). Moroccan saw the greatest benefit from this translation method, with F1 increases of 26.7% for EN and 11.4% for MSA.
@@ 6 Discussion
Our motivation for this study was to create a practical and reliable multilingual EQA benchmark dataset, without the need to start from scratch or resort to creating synthetic data. Given our languages of interest, repurposing the Belebele MCQA dataset seemed like a feasible solution. However, answering the research question on how feasible this approach is involves considering the various insights and findings presented in this paper.
We found that while repurposing is possible, our exploratory study showed that it required significant manual effort. Contrary to initial assumptions, we discovered that the original answers from the multiple-choice question pairs were not suitable for automated approaches. In fact, the only multiple-choice types in Table 1 that resemble non-complex SQUAD-style QA pairs are (a) Exact.
Rather than offering a scalable solution here, instead our manual efforts and the subsequent insights serve to make this repurposing task much more feasible and potentially automated by other dataset developers. We can see that the repurposing process is much more feasible when the differences between the QA styles are fully understood. We found that, through an effective filtering process, we could easily eliminate many QA pairs that did not lend themselves easily to EQA. While this resulted in reducing the dataset size by more than half, this leaves open new avenues for examining how those ‘unsuitable’ QA pairs could be reformulated to be useful in EQA. Repurposing a multilingual dataset like Belebele means that its parallelism could be exploited through a bootstrapping approach. We have learned that once one language dataset is manually annotated and verified, the spans could be relatively easily projected to a new language from the wider dataset provided a translation system is available (see Section 4.5). Likewise, few changes were required when adapting our annotation guidelines from EN to MSA, suggesting that this adaptability may hold for other languages. Finally, our experimental results confirm that our approach produced a useful data set that can be used for benchmarking cross-lingual EQA for all Belebele languages.
@@ 7 Conclusion and Future Work
We repurposed the MCQA Belebele dataset to create a new EQA dataset for English and MSA, with cross-lingual benchmark results for dialectal Arabic QA. [^10] https://github.com/mbzuai-nlp/MultiChoice2ExtractiveQA We also provided guidelines for practitioners exploring QA in other Belebele languages. Our work, including negative insights, aims to enhance understanding of the challenges and opportunities in reformulating NLP tasks.
For future work, we plan to support repurposing Belebele-EQA for other languages, potentially automating the process. The dataset offers a resource for evaluating such methods. We also aim to use it for fine-tuning general-purpose Arabic QA models and improving EQA evaluation metrics, focusing on span length.
@@
We would like to point to some limitations of our study. First, while we are addressing a general problem, in our study, out of all the other Belebele under-resourced languages, we only focused on cross-lingual evaluation of five Arabic dialects. While we believe that our findings are more generally applicable, we would like to note that each language and dataset may present unique challenges and opportunities. Second, our repurposing was based on a single dataset, Belebele, and for a single task (extractive question answering in the style of machine reading comprehension), and we note that other tasks may require some specific considerations.
@@
It is important to acknowledge the possibility of biases within our dataset due to subjective human judgments, or due to the influence of the original task’s data (e.g. a narrow cultural representation across the questions).
@@ Appendix A Processing IAA Spans for Visualization
Appendix B PrimeQA Experimental Details
The experiments discusses in Section 5.2 were conducted on an Apple M1 Max MacBook Pro with 32GBs of memory. Since we used the aforementioned models in inference mode, each experiment with 329 questions (E.g. the combination of En Passage and En Questions) took approximately 30 seconds to complete. Note that additional time may be required to download each model for the first time. The 56 core experiments (without translation) in Table 3 (2 Passages x 7 question types x 2 configurations x 2 normalization schemes) are expected to require 28 minutes of running.
Appendix C Arabic Examples of MCQ-Span Types
@@
Appendix D Guidelines for English Answer Span Annotation
See pages 2-6 of BelebeleQA-Annotation-Guidelines_EN_Paper_Appendix.pdf
Appendix E Guidelines for Modern Standard Arabic Answer Span Annotation